# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt

### v1

@@ Abstract
We investigate the robustness and security issues from a novel and practical setting: a group of malicious clients has impacted the model during training by disguising their identities and acting as benign clients, and only revealing their adversary position after the training to conduct transferable adversarial attacks with their data, which is usually a subset of the data that FL system is trained with. Our aim is to offer a full understanding of the challenges the FL system faces in this practical setting across a spectrum of configurations. We notice that such an attack is possible, but the federated model is more robust compared with its centralized counterpart when the accuracy on clean images is comparable. Through our study, we hypothesized the robustness is from two factors: the decentralized training on distributed data and the averaging operation. We provide evidence from both the perspective of empirical experiments and theoretical analysis. Our work has implications for understanding the robustness of federated learning systems and poses a practical question for federated learning applications.
@@ 1 Introduction
The proliferation of mobile devices leads to a surge in user-end data collection. If utilized effectively, this sensitive data can enhance intelligent systems. Federated Learning (FL) offers a decentralized learning solution, training models via local updates and parameter aggregation mcmahan2017communication . The FL system, comprising a federation of clients and a data-less centralized server, circulates the global model amongst a random subset of participants. Participants perform local updates and submit them to the server for aggregation. The system’s design leverages private data while preserving privacy.
@@
Although FL safeguards privacy, it is still susceptible to attacks such as data poisoning huang2011adversarial , model poisoning bhagoji2019analyzing ; bagdasaryan2020backdoor , free-riders attack lin2019free , and reconstruction attacks geiping2020inverting ; zhu2019deep . It is also vulnerable to adversarial attacks during inference biggio2013evasion ; szegedy2013intriguing , including adversarial examples designed to deceive the model zizzo2020fat . Research on robust FL methods against adversarial examples has primarily focused on a white-box setting where attackers have full model access zhou2020adversarial ; reisizadeh2020robust ; hong2021federated . However, real-world FL applications, like Gboard hard2018federated , usually restrict such access.
@@
We observe a distinct FL security challenge: malicious clients may pose as benign contributors, only revealing their adversarial intent post-training. These clients gain access to a subset of the training data, enabling potential adversarial attacks. Current FL applications lack effective mechanisms to eliminate such hostile participants hard2018federated , even if selection mechanisms exist, such as Krum fang2020local ; li2020learning ; bagdasaryan2020backdoor , as attackers do not exhibit hostile behavior during training. After obtaining data, an attacker could train a substitute model for transfer-based black-box attacks.
@@
In this paper, we pioneer an exploration of this practical perspective of FL robustness. Stemming from the above scenarios, we propose a simple yet practical assumption: the attacker possesses a limited amount of the users’ data but no knowledge about the target model or the full training set. To assess current FL system robustness and guide future research in this regard, we investigate the adversarial transferability under a spectrum of practical FL settings.
@@
We first evaluate the transferability of adversarial examples generated from different source models to attack a federated-trained model. Then a comprehensive evaluation of practical configurations is conducted to assess the feasibility of our attack. We further investigate two properties of FL, namely the decentralized training and the averaging operation, and their correlation with federated robustness. To provide a comprehensive evaluation in a practical aspect, we take into account the attack timing, the architecture and different aggregation methods in our experiments. We have the following findings:
@@
  • We investigate the transferability of adversarial examples generated from models trained by various number of users’ data. We observe that without any elaborated techniques such as dataset synthesis papernot2017practical or attention wu2020boosting , a regularly trained source model with only limited users’ data can perform the transfer attack. With ResNet50 on CIFAR10, we achieve a transfer rate of 70% and 80% with only 5% and 7% of users with augmentations and further improve this number to 81% and 85% with auto attack croce2020reliable .
  • We investigate two intrinsic properties of the FL, namely the property of distributed training and the averaging operation and discover that both heterogeneity and dispersion degree of the decentralized data as well as the averaging operations can significantly decrease the transfer rate of transfer-based black-box attack.
@@ 2.1 Adversarial Robustness
The adversarial robustness of a model is usually defined as the model’s ability to predict consistently in the presence of small changes in the input. Intuitively, if the changes to the image are so tiny that they are imperceptible to humans, these perturbation will not alter the prediction. Formally, given a well trained classifier $f$ and image-label pairs $(x,y)$ on which the model correctly classifies $f(x)=y$, $f$ is defined to be $\epsilon$-robust with respect to distance metric $d(\dot{,}\dot{)}$ if
@@
where $\alpha$ denotes the function evaluating prediction accuracy. In the case of classification, $\alpha(f(x;\theta),y)$ yields 1 if the prediction $f(x;\theta)$ equals ground-truth label $y$, 0 otherwise. The distance metric $d(\dot{,}\dot{)}$ is usually approximated by $L_{0}$, $L_{2}$ or $L_{\infty}$ to measure the visual dissimilarity of original image $x$ and the perturbed image $x^{\prime}$. Despite the change to the input is small, the community have found a class of methods that can easily manipulate model’s predictions by introducing visually imperceptible perturbations in images szegedy2013intriguing ; goodfellow2014explaining ; moosavi2016deepfool . From the optimization standpoint, it is achieved by maximizing the loss of the model on the input (madry2017towards, ):
@@
where $l(\cdot,\cdot)$ denotes the loss function (e.g. cross entropy loss) for training the model $f$ parameterized by $\theta$. While these attack methods are powerful, they usually require some degrees of knowledge about the target model $f$ (e.g. the gradient). Arguably, for many real-world settings, such knowledge is not available, and we can only expect less availability of such knowledge on FL applications trained and deployed by service providers. On the other hand, the hostile attacker having access to some but limited amount of users’ data is a much more realistic scenario. Thus, we propose the following assumption for practical attack in FL: given the data of $n$ malicious users $D_{m}=\bigcup_{i=1}^{n}D_{i}$ where $D_{i}=\{(x_{k},y_{k})|k=1,\cdots,m_{i}\}^{(i)}$ contains $m_{i}$ data points, we aim to acquire a transferable perturbation $\delta$ by maximizing the same objective as in Equation 3 but with a substitute model $f^{\prime}$ parameterized by $\theta^{\prime}$ trained by $D_{m}$:
$\delta=\mathop{\arg\min}_{\delta}\ l(f^{\prime}(x+\delta;\theta^{\prime}),y)\;\;\textnormal{s.t.}\;\;d(x+\delta,x)<\epsilon$ (4)
@@ 2.2 The Security and Robustness of Federated Learning
Poisoning and Backdoor Attack. Poisoning attacks biggio2012poisoning ; fang2020local aim to disrupt the global model by injecting malicious data or manipulating local model parameters on compromised devices. In contrast, backdoor attacks bagdasaryan2020backdoor ; sun2019can infuse a malicious task into the existing model without impacting its primary task accuracy, including model replacement chen2017targeted , label-flipping fung2018mitigating , and the use of GANs xie2019dba . Others have effectively incorporated weight boosting and stealth into the adversarial objective bhagoji2019analyzing . Defenses against these attacks often involve anomaly detection methods, such as Byzantine-tolerant aggregation shejwalkar2021manipulating (e.g., Krum, MultiKrum blanchard2017machine , Bulyan guerraoui2018hidden , Trimmed-mean and Median yin2018byzantine ), focusing on the geometric distance between hostile and benign gradients. Additional strategies include clustering shen2016auror , Error Rate and Loss based Rejection fang2020local , and learning a separate model to detect malicious updates li2020learning ; kieu2019outlier . The robustness and attack performance of backdoor attacks is significantly influenced by FL data heterogeneity zawad2021curse .
@@
Transfer Attack. Transfer-based adversarial attacks employ the full training set of the target model to train a substitute model zhou2018transferable , a challenging condition to meet in practice, especially in FL where data privacy is paramount. Another line of inquiry delves into the mechanisms of black-box attacks, which exploit the high transferability of adversarial examples even between different model architectures szegedy2013intriguing ; goodfellow2014explaining . This transferability is partially attributed to the similarity between source and target models goodfellow2014explaining ; liu2016delving , as adversarial perturbations align closely with a model’s weight vectors, and different models learn similar decision boundaries for the same task. tramer2017space found that adversarial examples span a large, contiguous subspace, facilitating transferability. Meanwhile, ilyas2019adversarial posits that adversarial perturbations are non-robust features captured by the model, and waseda2021closer utilizes this theory to explain differing mistakes in transfer attacks. Additionally, demontis2019adversarial reveals that similar gradients in source and target models and lower variance in loss landscapes increase transfer attack probability. We underscore our setting’s distinctiveness and importance compared to others.
@@
Key Difference 1: Different from query-based or transfer-based black-box attack, we assume the malicious clients possess the data themselves, impacting the target model during training and attack during inference. We also present a comparison of our attack setting and the query-based attack in Appendix A. Note that our attack setting doesn’t contradict with the query-based attack. In fact, we can perform with both if the FL system allows a certain number of queries, which we leave to future works to explore.
@@
Significance: Besides the potential data leakage by malicious participants, we also emphasize that despite, ideally, each participant having access to the global model, in real-world applications (e.g. Gboard hard2018federated ), the infrastructure provider will impose additional protection such as encryption or encapsulation over the local training. For instance, Gboard from Google provides the next-word prediction with FL, which requires users to install an app to participate. For an adversary, it’s impractical to obtain the global model without breaking or hacking the app or hijacking and decrypting the communication, despite all the things happening "locally". We believe this is much more difficult and laborsome than our setting which significantly boosts the transferability by simply acting as a benign.
@@ 3 Investigation Setup and Research Goals
With the possibility that malicious client disguise to be a benign during training and exploit the obtained data for adversarial attacks, as discussed in Section 1, it is of necessity to investigate the robustness of federated model against transfer-based attack, i.e. transferability of adversarial examples against FL. More importantly, we want to understand where do the robustness of FL comes from and how the core components of FL effects it. To this end, we put forward the following research goals:
@@ 3.1 Experiment Setup
Threat Model. Follwing zizzo2020fat , we use PGD madry2017towards with 10 iterations, no random restart, and an epsilon of 8 / 255 over $L_{\infty}$ norm on CIFAR10.
Settings. We first build up the basic FL setting. We split the datatset into 100 partitions of equal size in an iid fashion. We adopt two models for the experiments: CNN from mcmahan2017communication since it is commonly used in the FL literature and the widely used ResNet50he2016deep which represents a more realistic evaluation. We conduct training in three paradigms: the centralized model, federated model and the source model with limited number of clients’ data. For federated model, we use SGD without momentum, weight decay of 1e-3, learning rate of 0.1, local batch size of 50 following acar2021federated . We train locally 5 epochs on ResNet50 and 1 epoch on CNN. For centralized and source model training, we leverage SDG with momentum of 0.9, weight decay of 1e-3, learning rate of 0.01 and batch size of 64. For adversarial training, we use the same setting as centralized and leverage PGD to perform the adversarial training. We refer to zizzo2020fat for the detail of adversarial training. All experiments are conducted with one RTX 2080Ti GPU.
Metrics. We report Accuracy (Acc) and adversarial accuracy (Adv.Acc) for the performance and the robustness white-box attack, and for adversarial transferability, we report transfer accuracy (T.Acc) and transfer success rate (T.Rate) as detailed in 3.2.
@@ 3.2 Adversarial Transferability in Federated Learning
To define the transferability of adversarial examples, we first introduce the definition of the source model, target model and adversarial example. The source model is the substitute model used to generate adversarial examples while the target model is the target aimed to attack. Given the validation set $x=\{(x_{i},y_{i})\}$, source model $f^{\prime}$, target model $f$ and adversarial perturbation function $adv(\cdot,\cdot)$ (e.g. PGD), we first define the following sets: $s1=\{x_{i}|f^{\prime}(x_{i})=y_{i}\}$, $s2=\{x_{i}|f^{\prime}(adv(x_{i},f^{\prime}))\neq y_{i}\}$, $s3=\{x_{i}|f(x_{i})=y_{i}\}$, $s4=\{x_{i}|f(adv(x_{i},f^{\prime}))\neq y_{i}\}$. Adversarial examples are defined to be those samples that are originally correctly classified by model $f^{\prime}$ but are mis-classified when adversarial perturbation is added, i.e., $s1\cap s2$. Adversarial transferability against the target model refers to the ability of an adversarial examples generated from the source model to attack the target model (becomes an adversarial example of target model). We define transfer rate (T.Rate) and transfer accuracy (T.Acc) to measure the adversarial transferability:

### v2

# Towards Understanding Adversarial Transferability in Federated Learning

@@ Abstract
We investigate a specific security risk in FL: a group of malicious clients has impacted the model during training by disguising their identities and acting as benign clients but later switching to an adversarial role. They use their data, which was part of the training set, to train a substitute model and conduct transferable adversarial attacks against the federated model. This type of attack is subtle and hard to detect because these clients initially appear to be benign.
The key question we address is: How robust is the FL system to such covert attacks, especially compared to traditional centralized learning systems? We empirically show that the proposed attack imposes a high security risk to current FL systems. By using only 3% of the client’s data, we achieve the highest attack rate of over 80%. To further offer a full understanding of the challenges the FL system faces in transferable attacks, we provide a comprehensive analysis over the transfer robustness of FL across a spectrum of configurations. Surprisingly, FL systems show a higher level of robustness than their centralized counterparts, especially when both systems are equally good at handling regular, non-malicious data.
We attribute this increased robustness to two main factors: 1) Decentralized Data Training: Each client trains the model on its own data, reducing the overall impact of any single malicious client. 2) Model Update Averaging: The updates from each client are averaged together, further diluting any malicious alterations. Both practical experiments and theoretical analysis support our conclusions. This research not only sheds light on the resilience of FL systems against hidden attacks but also raises important considerations for their future application and development.
@@ 1 Introduction
Although Federated Learning (FL) provides a promising solution to collaboratively train models without exchanging data, especially in privacy-concerned areas Li et al. (2022), it is still susceptible to attacks such as data poisoning Huang et al. (2011), model poisoning Bhagoji et al. (2019); Bagdasaryan et al. (2020); Huang et al. (2023), free-riders attack Lin et al. (2019), and reconstruction attacks Geiping et al. (2020); Zhu et al. (2019). It is also vulnerable to adversarial attacks during inference Biggio et al. (2013); Szegedy et al. (2013), including adversarial examples designed to deceive the model Zizzo et al. (2020). Research on robust FL methods against adversarial examples has primarily focused on a white-box setting where attackers have full model access Zhou et al. (2020); Reisizadeh et al. (2020a); Hong et al. (2021); Qiao et al. (2024). However, real-world FL applications, like Gboard Hard et al. (2018), usually restrict such access.
@@
We observe a distinct FL security challenge: malicious clients may pose as benign contributors, only revealing their adversarial intent post-training. This setting raises a new security challenge because these clients have access to a subset of the training data, potentially leading to a better surrogate model for transferable attacks. Current FL applications lack effective mechanisms to eliminate such hostile participants Hard et al. (2018), even if selection mechanisms exist, such as Krum Fang et al. (2020); Li et al. (2020a); Bagdasaryan et al. (2020), as attackers do not exhibit hostile behavior during training. After obtaining data, an attacker could train a surrogate model for transfer-based black-box attacks.
@@
In this paper, we pioneer an exploration of this practical perspective of FL robustness. Stemming from the above scenarios, we propose a simple yet practical assumption: the attacker possesses a limited amount of the users’ data but no knowledge about the target model or the full training set. To assess current FL system robustness and guide future research in this regard, we investigate the adversarial transferability under a spectrum of practical FL settings. Adversarial transferability refers to the ability of adversarial examples generated from the source model to successfully attack the target model, which measures the amount of threat white-box attack poses to the model and system.
@@
We first evaluate the transferability of adversarial examples generated from different source models to attack a federated-trained model. Then a comprehensive evaluation of practical configurations is conducted to assess the feasibility of our attack. We further investigate two properties of FL: decentralized training and the averaging operation and their correlation with federated robustness. To provide a comprehensive evaluation in a practical aspect, we consider the attack timing, the architecture, and different aggregation methods in our experiments. We have the following findings:
@@
  • We investigate the transferability of adversarial examples generated from models trained by various numbers of users’ data. We observe that without any elaborated techniques such as dataset synthesis Papernot et al. (2017) or attention Wu et al. (2020b), a regularly trained source model with only limited users’ data can perform the transfer attack. With ResNet50 on CIFAR10, we achieve a transfer rate of 70% and 80% with only 5% and 7% of users with augmentations and further improve this number to 81% and 85% with AutoAttack Croce & Hein (2020).
  • We investigate two intrinsic properties of the FL: the property of distributed training and the averaging operation and discover that both heterogeneity and dispersion degree of the decentralized data as well as the averaging operations can significantly decrease the transfer rate of transfer-based black-box attack.
@@ 2.1 Adversarial Robustness
The adversarial robustness of a model is usually defined as the model’s ability to predict consistently in the presence of small changes in the input. Intuitively, if the changes to the image are so tiny that they are imperceptible to humans, these perturbations will not alter the prediction. Formally, given a well trained classifier $f$ and image-label pairs $(x,y)$ on which the model correctly classifies $f(x)=y$, $f$ is defined to be $\epsilon$-robust with respect to distance metric $d(\dot{,}\dot{)}$ if
@@
where $\alpha$ denotes the function evaluating prediction accuracy. In the case of classification, $\alpha(f(x;\theta),y)$ yields 1 if the prediction $f(x;\theta)$ equals ground-truth label $y$, 0 otherwise. The distance metric $d(\dot{,}\dot{)}$ is usually approximated by $L_{0}$, $L_{2}$ or $L_{\infty}$ to measure the visual dissimilarity of original image $x$ and the perturbed image $x^{\prime}$. Despite the change to the input is small, the community have found a class of methods that can easily manipulate model’s predictions by introducing visually imperceptible perturbations in images Szegedy et al. (2013); Goodfellow et al. (2014); Moosavi-Dezfooli et al. (2016). From the optimization standpoint, it is achieved by maximizing the loss of the model on the input Madry et al. (2017):
@@
where $l(\cdot,\cdot)$ denotes the loss function (e.g., cross-entropy loss) for training the model $f$ parameterized by $\theta$. While these attack methods are powerful, they usually require some degrees of knowledge about the target model $f$ (e.g., the gradient). Arguably, for many real-world settings, such knowledge is not available, and we can only expect less availability of such knowledge on FL applications trained and deployed by service providers. On the other hand, the hostile attacker having access to some but limited amount of users’ data is a much more realistic scenario. Thus, we propose the following assumption for practical attack in FL: given the data of $n$ malicious users $D_{m}=\bigcup_{i=1}^{n}D_{i}$ where $D_{i}=\{(x_{k},y_{k})|k=1,\cdots,m_{i}\}^{(i)}$ contains $m_{i}$ data points, we aim to acquire a transferable perturbation $\delta$ by maximizing the same objective as in Equation 3 but with a surrogate model $f^{\prime}$ parameterized by $\theta^{\prime}$ trained by $D_{m}$:
$\delta=\mathop{\arg\max}_{\delta}\ l(f^{\prime}(x+\delta;\theta^{\prime}),y)\;\;\textnormal{s.t.}\;\;d(x+\delta,x)<\epsilon$ (4)
@@ 2.2 The Security and Robustness of Federated Learning
Poisoning attacks Biggio et al. (2012); Fang et al. (2020) aim to disrupt the global model by injecting malicious data or manipulating local model parameters on compromised devices. In contrast, backdoor attacks Bagdasaryan et al. (2020); Sun et al. (2019); Wang et al. (2023) infuse a malicious task into the existing model without impacting its primary task accuracy Zhang et al. (2023); Huang et al. (2024), including model replacement Chen et al. (2017), label-flipping Fung et al. (2018), fixed-trigger backdoor Dai & Li (2023); Liu et al. (2024) and trigger-optimization Huang (2020); Li et al. (2023a); Nguyen et al. (2024). Defenses against these attacks often involve anomaly detection methods, such as Byzantine-tolerant aggregation Shejwalkar & Houmansadr (2021) (e.g., Krum, MultiKrum Blanchard et al. (2017), Bulyan Guerraoui et al. (2018), Trimmed-mean and Median Yin et al. (2018)), focusing on the geometric distance between hostile and benign gradients. More advanced defenses take detection, aggregation, detection, and differential privacy into consideration Huang et al. (2023); Nguyen et al. (2022). The robustness and attack performance of backdoor attacks is significantly influenced by FL data heterogeneity Zawad et al. (2021).
Transfer Attack.
@@
Transfer-based adversarial attacks employ the full training set of the target model to train a surrogate model Zhou et al. (2018), a challenging condition to meet in practice, especially in FL where data privacy is paramount. Another line of inquiry delves into the mechanisms of black-box attacks, which exploit the high transferability of adversarial examples even between different model architectures Szegedy et al. (2013); Goodfellow et al. (2014). This transferability is partially attributed to the similarity between source and target models Goodfellow et al. (2014); Liu et al. (2016); Li et al. (2023b), as adversarial perturbations align closely with a model’s weight vectors, and different models learn similar decision boundaries for the same task. Tramèr et al. (2017) found that adversarial examples span a large, contiguous subspace, facilitating transferability. Meanwhile, Ilyas et al. (2019) posits that adversarial perturbations are non-robust features captured by the model, and Waseda et al. (2023) utilizes this theory to explain differing mistakes in transfer attacks. Additionally, Demontis et al. (2019); Zhang et al. (2024) reveals that similar gradients in source and target models and lower variance in loss landscapes increase transfer attack probability. Despite transfer attacks being more realistic and practical, not much attention has been focused on this aspect of the safety and robustness of FL. We take the initiative to investigate this area and underscore our setting’s distinctiveness and importance compared to others.
@@
Key Difference 1: Different from query-based or transfer-based black-box attack, we assume the malicious clients possess the data themselves, impacting the target model during training and attack during inference. We also present a comparison of our attack setting and the query-based attack in Section 4.3. Note that our attack setting doesn’t contradict the query-based attack. In fact, we can perform with both if the FL system allows a certain number of queries, which we leave to future works to explore.
@@
Significance: Besides the potential data leakage by malicious participants, we also emphasize that despite, ideally, each participant having access to the global model, in real-world applications (e.g. Gboard Hard et al. (2018)), the infrastructure provider will impose additional protection such as encryption or encapsulation over the local training. For instance, Google’s Gboard provides next-word prediction with FL, which requires users to install an app to participate. For an adversary, it’s impractical to obtain the global model without breaking or hacking the app or hijacking and decrypting the communication, despite all the things happening "locally". We believe this is much more difficult and laborsome than our setting which significantly boosts the transferability by simply acting as a benign.
@@ 3.1 Experiment Setup
Threat Model. Following (Zizzo et al., 2020), we use PGD (Madry et al., 2017) with 10 iterations, no random restart, and an epsilon of 8 / 255 over $L_{\infty}$ norm on CIFAR10. For experiments on ImageNet200, we use PGD (Madry et al., 2017) of the same setup but with an epsilon of 2 / 255.
Settings. We first build up the basic FL setting. We split the datatset into 100 partitions of equal size in an iid fashion. We adopt two models for the experiments: CNN from (McMahan et al., 2017) since it is commonly used in the FL literature and the widely used ResNet50 (He et al., 2016) which represents a more realistic evaluation. We conduct training in three paradigms: the centralized model, the federated model and the source model with a limited number of clients’ data. For the federated model, we use SGD without momentum, weight decay of 1e-3, learning rate of 0.1, and local batch size of 50 following (Acar et al., 2021). We follow the cross-device setting and use a subset of clients in each round of training. We use a 10% as default where not specified. We train locally 5 epochs on ResNet50 and 1 epoch on CNN. For centralized and source model training, we leverage SGD with a momentum of 0.9, weight decay of 1e-3, a learning rate of 0.01 and batch size of 64. For adversarial training, we use the same setting as centralized and leverage PGD to perform the adversarial training. We refer to (Zizzo et al., 2020) for the details of adversarial training. All experiments are conducted with one RTX 2080Ti GPU.
Metrics. We report Accuracy (Acc) and adversarial accuracy (Adv.Acc) for the performance and the robustness of white-box attack, and for adversarial transferability, we report transfer accuracy (T.Acc) and transfer success rate (T.Rate) as detailed in 3.2.
@@ 3.2 Adversarial Transferability in Federated Learning
To define the transferability of adversarial examples, we first introduce the definition of the source model, target model and adversarial example. The source model is the surrogate model used to generate adversarial examples while the target model is the target aimed to attack. Given the validation set $x=\{(x_{i},y_{i})\}$, source model $f^{\prime}$, target model $f$ and adversarial perturbation function $adv(\cdot,\cdot)$ (e.g., PGD), we first define the following sets:
$\large s1$ =$=\{x_{i}|f^{\prime}(x_{i})=y_{i}\},$
$s2$ =$=\{x_{i}|f^{\prime}(adv(x_{i},f^{\prime}))\neq y_{i}\},$
$s3$ =$=\{x_{i}|f(x_{i})=y_{i}\},$
$s4$ =$=\{x_{i}|f(adv(x_{i},f^{\prime}))\neq y_{i}\}$